An urn contains 27 red marbles, 25 blue marbles, and 43 yellow marbles. One marble is to be chosen from the urn without looking. What is the probability of choosing a red marble?

Your answer should be to rounded to 4 decimal places.
Answer #1

Total marbles = 27 + 25 + 43 = 95

P(Red marble) = n(Red marble) / Total marbles

= 27 / 95

= 0.2842